OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

slope (3+6)/(3+6)=1 y=x x-y=0

OpenStudy (alexandervonhumboldt2):

x-y=0

OpenStudy (anonymous):

thnx